Planned spending is as follows: C = 600 + 0.7(Y – T) – 200r, IP = 200 – 500r, G = 200, NX = 10 and T = 150. Which of the following also states PAE?

 

A. [905 – 700r] + 0.3Y

B. [905 – 700r]

C. [910 – 700r]

D. [905 – 700r] + 0.7Y

1
Expert's answer
2021-11-26T12:12:47-0500

PAE (planned aggregate expenditure) is a sum of consumption spending, investment spending, government spending, and net exports.

PAE=C+IP+G+NX

PAE=600+0.7(Y−150)−200r+200−500r+200+10

PAE=600+0.7Y−105−200r+200−500r+200+10

PAE=905−700r+0.7Y

PAE=[905−700r]+0.7Y

PAE=C+IP+G+NXPAE=600+0.7Y-150-200r+200-500r+200+10PAE=600+0.7Y-105-200r+200-500r+200+10PAE=905-700r+0.7YPAE=905-700r+0.7Y

The correct option is D.